Can i get help with finding the length of the sides of the original square
If the sides of a square are increased by 5 m, the area becomes 100 sq. meters. Find the length of the sides of the original square.

Answer by lwsshak3(11628)

Can i get help with finding the length of the sides of the original square
If the sides of a square are increased by 5 m, the area becomes 100 sq. meters. Find the length of the sides of the original square.
**
let x=side of original square
x+5 = side of square with area=100 sq. meters
(x+5)^2=100
take sqrt of both sides
x+5=10
x=5
ans
Length of side of original square=5m